# Unfolding the kagome lattice to improve understanding of ARPES in CoSn

### v1

@@ VI Conclusion
We have shown how bands can be separated according to their parity through polarization analysis adapted to the kagome lattice and how systematic variation of the ARPES band intensities in neighboring Brillouin zones can be captured by relatively simple unfolded DFT calculations. This provides additional tools to characterize the nature of a band beyond its dispersion. In the non-magnetic kagome compound CoSn, isolating each dispersion suggests that the bands are not identically renormalized, as usually assumed, but that the band crossing the Fermi level is renormalized by a factor of 2, pulling up all other bands by as much as 0.2 eV. This manifestation of electronic correlations appears quite different from what was observed in other multi-orbital systems, such as iron pnictides Yi et al. (2017), where the renormalization rather seems to depend on orbital character. This could be interesting to investigate further.
Following this idea, if many bands cross the Fermi level with different occupied bandwidths, one could expect a rather complex pattern of shifts compared to DFT calculations. In this case, it becomes crucial to have additional ways to identify the nature of a band to make sense of a comparison between ARPES and DFT. This is probably what happens in the magnetic FeSn, where many bands cross the Fermi level and the comparison with DFT calculations is indeed not straightforward at all anymore Ghimire and Mazin (2020). Symmetry-resolved unfolded calculations should help achieve a rational understanding of the ARPES measurement.
@@ Appendix B Extended zone scheme
Standard band unfolding techniques rely on a known relationship between the supercell and an assumed primitive cell Wang et al. (1998); Boykin and Klimeck (2005). Identifying the parent cell is often straightforward in systems with structural distortions (e.g., octahedral tilting Rubel et al. (2023)), magnetic order, or disordered alloys modeled via random atomic substitutions in a supercell Rubel et al. (2014). However, in some cases, superlattice effects are more subtle and difficult to disentangle. A prominent example is quasicrystals, which display sharp Bragg peaks despite lacking conventional long-range periodicity Bindi et al. (2009). The extended zone scheme provides a way to represent the electronic structure without assuming a specific primitive cell or Brillouin zone. This approach has previously been applied to aperiodic systems to extract a quasi-dispersion relation Niizeki and Akamatsu (1990); Mayo et al. (2020).
@@
Here, labels the atomic site positions, is the total number of sites, denotes the site amplitude, and is an atom-centered basis function. Due to the lack of translational symmetry, this state does not possess a well-defined crystal momentum. The spectral function in the extended zone scheme is obtained via the generalized structure factor Niizeki and Akamatsu (1990)
This result is analogous to the diffraction pattern of a quasicrystal, where serves as a site-dependent scattering amplitude Niizeki and Akamatsu (1990).
@@
The Dirac delta function effectively selects a single plane wave component, satisfying , that contributes to the spectral weight. Alternative formulations of the spectral function in the extended zone, aimed at analyzing ARPES measurements, have also been proposed in the literature Kosugi et al. (2017); Mayo et al. (2020).
@@
Figure 8(a) shows a slice through the Fermi surface of CoSn calculated using the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996, 1996) and presented in the extended zone scheme. A circular pattern appears in the first Brillouin zone but is absent in the adjacent periodic image, reappearing only in the third shell. This doubling of periodicity in reciprocal space suggests a Brillouin zone that is twice as large, which corresponds to a perceived primitive cell that is twice as small in real space, as illustrated in Fig. 8(b). This result supports the validity of the unfolding method and confirms the appropriateness of the chosen primitive cell.
@@ Appendix C ARPES matrix elements
Comprehensive treatments of ARPES principles can be found in textbooks Hüfner (2013); Tejeda and Malterre (2024) and papers Damascelli et al. (2003); Moser (2017); Yen et al. (2024). Applying Fermi’s golden rule, the photoemission intensity will depend on matrix elements of the form:
where describes the interaction between electron and photon, and are initial and final states. These are usually -particle states, but for our discussion, only the one-electron matrix element will matter. is usually developed as a perturbation in the dipolar approximation , although this is not sufficient for some cases (especially when the Hamiltonian is non-local Hwang et al. (2011)). The final state is usually taken as a plane wave of wave-vector , assuming the photoelectron is free to travel to the detector. Corrections to this approximation are often important to properly describe the dependence on photon energy or polarization Kern et al. (2023). Unfortunately, this often requires sophisticated calculations, which miss a simple connection with the properties of the initial state.
@@
To calculate the ARPES intensity, we follow Nishimoto et al., who derived Eq. (2) in the case of graphite Nishimoto et al. (1996). A similar derivation can be found elsewhere Moser (2017); Yen et al. (2024). The term in brackets essentially depends on the shape of the orbital at site . They are all the same in our case and we call =, where is a lattice site. Recalling the final state is a Bloch function of wave vector , we have .

### v2

@@ VI Conclusion
We have shown how bands can be separated according to their parity through polarization analysis adapted to the kagome lattice and how systematic variation of the ARPES band intensities in neighboring Brillouin zones can be captured by relatively simple unfolded DFT calculations. This provides additional tools to characterize the nature of a band beyond its dispersion. In the non-magnetic kagome compound CoSn, isolating each dispersion suggests that the bands are not identically renormalized, as usually assumed, but that the band crossing the Fermi level is renormalized by a factor of 2, pulling up all other bands by as much as 0.2 eV. This manifestation of electronic correlations appears quite different from what was observed in other multi-orbital systems, such as iron pnictides [47], where the renormalization rather seems to depend on orbital character. This could be interesting to investigate further.
Following this idea, if many bands cross the Fermi level with different occupied bandwidths, one could expect a rather complex pattern of shifts compared to DFT calculations. In this case, it becomes crucial to have additional ways to identify the nature of a band to make sense of a comparison between ARPES and DFT. This is probably what happens in the magnetic FeSn, where many bands cross the Fermi level and the comparison with DFT calculations is indeed not straightforward at all anymore [46] and probably complicated by surface contributions [12, 13]. Symmetry-resolved unfolded calculations should help achieve a rational understanding of the ARPES measurement.
Acknowledgement
We acknowledge Victor Balédent and the MORPHEUS, MESURES PHYSIQUES and MICRO-NANO platforms of the Laboratoire de Physique des Solides, Université Paris-Saclay, for help with sample characterization and Alberto Zobelli for introducing us to ref. 30. We acknowledge SOLEIL for provision of synchrotron radiation facilities at the CASSIOPEE beamline (proposal 20221212 and 20230669), the MAX IV Laboratory for beamtime on the BLOCH beamline under proposal 20231202 and Craig Polley and Mats Leandersson for their help during these measurements. Research conducted at MAX IV, a Swedish national user facility, is supported by Vetenskapsrådet (Swedish Research Council, VR) under contract 2018-07152, Vinnova (Swedish Governmental Agency for Innovation Systems) under contract 2018-04969 and Formas under contract 2019-02496. O.R. acknowledges computing resources provided by the Digital Research Alliance of Canada and travel support from the Natural Sciences and Engineering Research Council of Canada under Discovery Grant RGPIN-2025-05413.
@@ Appendix B Extended zone scheme
Standard band unfolding techniques rely on a known relationship between the supercell and an assumed primitive cell [48, 49]. Identifying the parent cell is often straightforward in systems with structural distortions (e.g., octahedral tilting [43]), magnetic order, or disordered alloys modeled via random atomic substitutions in a supercell [28]. However, in some cases, superlattice effects are more subtle and difficult to disentangle. A prominent example is quasicrystals, which display sharp Bragg peaks despite lacking conventional long-range periodicity [50]. The extended zone scheme provides a way to represent the electronic structure without assuming a specific primitive cell or Brillouin zone. This approach has previously been applied to aperiodic systems to extract a quasi-dispersion relation [51, 52].
@@
Here, labels the atomic site positions, is the total number of sites, denotes the site amplitude, and is an atom-centered basis function. Due to the lack of translational symmetry, this state does not possess a well-defined crystal momentum. The spectral function in the extended zone scheme is obtained via the generalized structure factor [51]
This result is analogous to the diffraction pattern of a quasicrystal, where serves as a site-dependent scattering amplitude [51].
@@
The Dirac delta function effectively selects a single plane wave component, satisfying , that contributes to the spectral weight. Alternative formulations of the spectral function in the extended zone, aimed at analyzing ARPES measurements, have also been proposed in the literature [53, 52].
@@
Figure 8(a) shows a slice through the Fermi surface of CoSn calculated using the Vienna ab initio simulation package (VASP) [54, 55, 56] and presented in the extended zone scheme. Exchange-correlation effects were described within the generalized-gradient approximation of Perdew, Burke, and Ernzerhof (PBE) [34]. Projector-augmented-wave potentials (version 5.4) [57] were employed, with 15 and 4 valence electrons for Co and Sn, respectively. The cutoff energy for the plane-wave expansion was set to 271 eV, corresponding to the maximum ENMAX value among the atomic species in the POTCAR file. A -centered k-mesh was used to sample the BZ. A Gaussian smearing of width 0.01 eV was applied to facilitate BZ integration and ensure smooth convergence of the electronic self-consistent cycle. The electronic energy convergence threshold was set to eV. SOC was included.
A circular pattern appears in the first Brillouin zone but is absent in the adjacent periodic image, reappearing only in the third shell. This doubling of periodicity in reciprocal space suggests a Brillouin zone that is twice as large, which corresponds to a perceived primitive cell that is twice as small in real space, as illustrated in Fig. 8(b). This result supports the validity of the unfolding method and confirms the appropriateness of the chosen primitive cell.
@@ Appendix C ARPES matrix elements
Comprehensive treatments of ARPES principles can be found in textbooks [58, 59] and papers [17, 37, 60]. Applying Fermi’s golden rule, the photoemission intensity will depend on matrix elements of the form:
where describes the interaction between electron and photon, and are initial and final states. These are usually -particle states, but for our discussion, only the one-electron matrix element will matter. is usually developed as a perturbation in the dipolar approximation , although this is not sufficient for some cases (especially when the Hamiltonian is non-local [61]). The final state is usually taken as a plane wave of wave-vector , assuming the photoelectron is free to travel to the detector. Corrections to this approximation are often important to properly describe the dependence on photon energy or polarization [62]. Unfortunately, this often requires sophisticated calculations, which miss a simple connection with the properties of the initial state.
@@
To calculate the ARPES intensity, we follow Nishimoto et al., who derived Eq. (2) in the case of graphite [20]. A similar derivation can be found elsewhere [37, 60]. The term in brackets essentially depends on the shape of the orbital at site . They are all the same in our case and we call =, where is a lattice site. Recalling the final state is a Bloch function of wave vector , we have .